OpenStudy (anonymous):

Write an equation in slope-intercept, point-slope, or standard form for th line with the given information. 1.) passes through (-1,4) and (-5,2) 2.) slope 2, y-intercept -4 3.) has an x-intercept of 6 and a y-intercept of 3 4.) passes through (1,2) with slope of -5/3

OpenStudy (anonymous):

2) y = 2x - 4

OpenStudy (anonymous):

all you have to do for this is find the slope and find the point where x = 0

OpenStudy (anonymous):

not sure how to do this though

OpenStudy (anonymous):

slope = y2 - y1/ x2 - x1 so for 1) slope = 2-4/-5-(-1) slope = 1/2 and then use any one of those two points

OpenStudy (anonymous):

so the slope intercept for 1) is 2-4/-5-(-1)

OpenStudy (anonymous):

ya, google your question, you should find lot of answers or similar questions